tiktokstar:

(05.02)An artist is designing a kite like the one show below. Calculate the area to determine how much material she will need to create the kite. A kite with two top triangles with height of 7 inches and a base 9.5 inches. The bottom two triangles share the base but have a height labeled 12 inches. 38 square inches 66.5 square inches 114 square inches 180.5 square inches

tiktokstar:

1 attachment
tiktokstar:

help plzzzzzz

jimthompson5910:

One diagonal is 7+12 = 19 inches The other diagonal is 9.5+9.5 = 19 inches In this case, both diagonals are the same length (though other times this won't be the case) To find the area of a kite, you multiply the length of the diagonals and then divide by 2 \(\large \text{Area} = \frac{a*b}{2}\) a,b are the two diagonal lengths. So you'll plug in a = 19 and b = 19.

tiktokstar:

so it is 19?

jimthompson5910:

\(\large \text{area} = \frac{a*b}{2} = \frac{19*19}{2} = ??\)

jimthompson5910:

The answer isn't 19

tiktokstar:

sooooooo what is it plz tell me

jimthompson5910:

what is 19 times 19 equal to?

tiktokstar:

361

tiktokstar:

right?

jimthompson5910:

Yes, now we cut that in half 361/2 is equal to what?

tiktokstar:

180.5

jimthompson5910:

Correct. That's the final answer

tiktokstar:

ok yay thanks

jimthompson5910:

No problem